OpenStudy (seattle12345):

Can someone please help with this?x^-1+10x^-2+25x^-3

OpenStudy (zehanz):

You mean you have this form: \[\frac{ 1 }{ x }+\frac{ 10 }{ x^2 }+\frac{ 25 }{ x^3 }\] What are you supposed to do with it? Write it as one fraction?

OpenStudy (anonymous):

(x^2/x^3)+(10x/x^3)+(25+x^3)

OpenStudy (anonymous):

(x^2+10x+25/x^3)

OpenStudy (anonymous):

\[(x+5)^2/x^3\]

OpenStudy (zehanz):

Well, if writing it as one fraction was what you were supposed to do, then madscientist101 has answered your question...

OpenStudy (anonymous):

I'm assuming it's simplification, however, ZeHanz is right, you need to tell us what to do with it.

OpenStudy (seattle12345):

Its factoring with negative exponents the answer is (x+5)^2/x^3 but not sure how to get there

hero (hero):

Bro, @madscientist pretty much posted the solution for you

OpenStudy (anonymous):